# SVGDreamer: Text Guided SVG Generation with Diffusion Model

### 03 Jan 2024 v2

@@ Abstract
Recently, text-guided scalable vector graphics (SVGs) synthesis has shown promise in domains such as iconography and sketch. However, existing text-to-SVG generation methods lack editability and struggle with visual quality and result diversity. To address these limitations, we propose a novel text-guided vector graphics synthesis method called SVGDreamer. SVGDreamer incorporates a semantic-driven image vectorization (SIVE) process that enables the decomposition of synthesis into foreground objects and background, thereby enhancing editability. Specifically, the SIVE process introduce attention-based primitive control and an attention-mask loss function for effective control and manipulation of individual elements. Additionally, we propose a Vectorized Particle-based Score Distillation (VPSD) approach to tackle the challenges of color over-saturation, vector primitives over-smoothing, and limited result diversity in existing text-to-SVG generation methods. Furthermore, on the basis of VPSD, we introduce Reward Feedback Learning (ReFL) to accelerate VPSD convergence and improve aesthetic appeal. Extensive experiments have been conducted to validate the effectiveness of SVGDreamer, demonstrating its superiority over baseline methods in terms of editability, visual quality, and diversity. The code and demo of SVGDreamer can be found at https://ximinng.github.io/SVGDreamer-project/.
@@ 1 Introduction
Scalable Vector Graphics (SVGs) represent visual concepts using geometric primitives such as Bezier curves, polygons, and lines. Due to their inherent nature, SVGs are highly suitable for visual design applications, such as posters and logos. Firstly, vector images are not constrained by resolution and can print results in any size without compromising clarity. Secondly, compared to raster images, vector images can maintain compact file sizes, making them more efficient for storage and transmission purposes. More importantly, vector images offer greater editability, allowing designers to easily select, modify, and compose elements. This attribute is particularly crucial in the design process, as it allows for seamless adjustments and creative exploration.
@@
In recent years, there has been a growing interest in general vector graphics generation. Various optimization-based methods [4, 29, 20, 40, 41, 34, 12, 48] have been proposed, building upon the differentiable rasterizer DiffVG [14]. These methods, such as CLIPDraw [4] and VectorFusion [12], differ primarily in their approach to supervision. Some works [4, 29, 20, 34, 40, 41] combine the CLIP model [24] with DiffVG [14], using CLIP as a source of supervision. More recently, the significantly progress achieved by Text-to-Image (T2I) diffusion models [21, 27, 25, 28, 37] has inspired the task of text-to-vector-graphics. Both VectorFusion [12] and DiffSketcher [48] attempted to utilize T2I diffusion models for supervision. These models make use of the high-quality raster images generated by T2I models as targets to optimize the parameters of vector images. Additionally, the priors embedded within T2I models can be distilled and applied in this task. As a result, models using T2I models for supervision generally achieve better performance compared to those using the CLIP model.
@@
Despite their impressive performance, existing T2I-based methods have certain limitations. Firstly, the vector images generated by these methods lack editability. Unlike the conventional approach of creating vector graphics, where individual elements are added one by one, T2I-based methods do not distinguish between different components during synthesis. As a result, the objects become entangled, making it challenging to edit or modify a single object independently. Secondly, there is still a large room for improvement in visual quality and diversity of the results generated by these methods. Both VectorFusion and DiffSketcher use data Augmentation extended Score Distillation Sampling (SDS) [23] to distill priors from the T2I models. However, it has been observed that SDS can lead to issues such as color over-saturation and over-smoothing, resulting in a lack of fine details in the generated vector images. Besides, SDS optimizes a set of control points in the vector graphic space to obtain the average state of the vector graphic corresponding to the text prompt in a mode-seeking manner [23]. This leads to a lack of diversity and detailed construction in the SDS-based approach [12, 48], along with absent text prompt objects.
@@
To address the aforementioned issues, we present a new model called SVGDreamer for text-guided vector graphics generation. Our primary objective is to produce vector graphics of superior quality that offer enhanced editability, visual appeal, and diversity. To ensure editability, we propose a semantic-driven image vectorization (SIVE) process. This approach incorporates an innovative attention-based primitive control strategy, which facilitates the decomposition of the synthesis process into foreground objects and background. To initialize the control points for each foreground object and background, we leverage a cross-attention map queried by text tokens. Furthermore, we introduce an attention-mask loss function, which optimizes the graphic elements hierarchically. The proposed SIVE process ensures the separation and editability of the individual elements, promoting effective control and manipulation of the resulting vector graphics.
@@
To improve the visual quality and diversity of the generated vector graphics, we introduce two new techniques: Vectorized Particle-based Score Distillation (VPSD) and Reward Feedback Learning (ReFL) for vector graphics refinement. Previous works in vector graphics synthesis [12, 48, 11] that utilized score distillation sampling [23] often encountered issues like color over-saturation, color over-smoothing of vector shapes, and limited diversity in the synthesis results [23, 48]. To address these challenges, we propose a novel method that enables richer color effects and fosters greater diversity in the synthesized vector graphics. We model a vector graphic as distributions of control points and colors, respectively, and apply VPSD to synthesize the graphics’ contents. Furthermore, to enhance the aesthetic appeal of the generated vector graphics, we introduce ReFL [49] to fine-tune the estimation network for the variational distribution. Through this refinement process, we achieve the final vector graphics that exhibit high editability, superior visual quality, and increased diversity. To validate the effectiveness of our proposed method, we perform extensive experiments to evaluate the model across multiple aspects. In summary, our contributions can be summarized as follows:
@@
  • We present the semantic-driven image vectorization (SIVE) method, which ensures that the generated vector objects are separate and flexible to edit. Additionally, we propose the vectorized particle-based score distillation (VPSD) loss and Reward Feedback Learning to guarantee that the generated vector graphics exhibit both exceptional visual quality and a wide range of diversity.
@@ 2.1 Vector Graphics Generation
Scalable Vector Graphics (SVGs) offer a declarative format for visual concepts expressed as primitives. One approach to creating SVG content is to use Sequence-To-Sequence (seq2seq) models to generate SVG scripts [5, 16, 1, 26, 43, 44, 46]. These methods heavily rely on dataset in vector form, which limits their generalization ability and their capacity to synthesize complex vector graphics. Instead of directly learning an SVG generation network, an alternative method of vector synthesis is to optimize towards a matching image during evaluation time.
Li et al. [14] introduce a differentiable rasterizer that bridges the vector graphics and raster image domains. While image generation methods that traditionally operate over vector graphics require a vector-based dataset, recent work has demonstrated the use of differentiable renderers to overcome this limitation. [31, 39, 26, 29, 18, 38, 36, 48]. Furthermore, recent advances in visual text embedding contrastive language-image pre-training model (CLIP) [24] have enabled a number of successful methods for synthesizing sketches, such as CLIPDraw[4], CLIP-CLOP [20], and CLIPasso [40]. A very recent work VectorFusion [12] and DiffSketcher [48] combine differentiable renderer with text-to-img diffusion model for vector graphics generation, resulting in promising results in fields such as iconography, pixel art, and sketch.
@@ 2.2 Text-to-Image Diffusion Model
Denoising diffusion probabilistic models (DDPMs) [32, 33, 8, 35], particularly those conditioned on text, have shown promising results in text-to-image synthesis. For example, Classifier-Free Guidance (CFG) [7] has improved visual quality and is widely used in large-scale text conditional diffusion model frameworks, including GLIDE [21], Stable Diffusion [27], DALL·E 2 [25], Imagen [28] and DeepFloyd IF [37]. The significant progress achieved by text-to-image diffusion models [21, 27, 25, 28] also promote the development of a series of text-guided tasks, such as text-to-3D [23]. In this work, we employ Stable Diffusion model to provide supervision for text-to-SVG generation.
@@ 3 Methodology
Figure 2: Overview of our proposed SVGDreamer. The method consists of two parts: semantic-driven image vectorization (SIVE) and SVG synthesis through VPSD optimization. The result obtained from SIVE can be used as input for VPSD, which helps in establishing a dependable initialization point.
@@ 3.2 Vectorized Particle-based Score Distillation
While vectorizing a rasterized diffusion sample is lossy, recent techniques [12, 48, 11] have identified the SDS loss [23] as beneficial for our task of generating vector graphics. To synthesize a vector image that matches a given text prompt $y$, they directly optimize the parameters $\theta=\{P_{i},C_{i}\}_{i=1}^{n}$ of the differentiable rasterizer $\mathcal{R}(\theta)$ via SDS loss. At each iteration, a differentiable rasterizer is used to render a raster image $\mathbf{x}=\mathcal{R}(\theta)$, which is augmented to obtain a $\mathbf{x}_{a}$. Then, the pretrained latent diffusion model (LDM) $\epsilon_{\phi}$ uses a VAE encoder [3] to encode $\mathbf{x}_{a}$ into a latent representation $\mathbf{z}=\mathcal{E}(\mathbf{\mathbf{x}_{a}})$, where $\mathbf{z}\in\mathbb{R}^{(H/f)\times(W/f)\times 4}$ and $f$ is the encoder downsample factor. Finally, the gradient of SDS is estimated by,
@@
Unfortunately, SDS-based methods often suffer from issues such as color over-saturation, vector graphic over-smoothing, and limited diversity in synthesis results [23, 12, 48, 11].
Figure 3: An illustration of the Vectorized Particle-based Score Distillation (VPSD).
Inspired by the principled variational score distillation framework [45], we propose vectorized particle-based score distillation (VPSD) to address the aforementioned issues. Instead of modeling vector graphics as a set of control points and corresponding colors like SDS, we model vector graphics as the distribution of control points and colors respectively. In principle, given a text prompt $y$, there exists a probabilistic distribution $\mu$ of all possible vector shapes representations. Under a vector representation parameterized by $\theta$, such a distribution can be modeled as a probabilistic density $\mu(\theta|y)$. Compared with SDS that optimizes for the single $\theta$, VPSD optimizes for the whole distribution $\mu$, from which we can sample $\theta$. Motivated by previous particle-based variational inference methods, we maintain $k$ groups of vector parameters $\{\theta\}_{i=1}^{k}$ as particles to estimate the distribution $\mu$, and $\theta(i)$ will be sampled from the optimal distribution $\mu^{\ast}$ if the optimization converges. This optimization can be realized through two score functions: one that approximates the optimal distribution with a noisy real image, and one that represents the current distribution with a noisy rendered image. The score function of noisy real images can be approximated by the pretrained diffusion model [27] $\mathbf{\epsilon}_{\phi}(\mathbf{z}_{t};y,t)$. The score function of noisy rendered images is estimated by another noise prediction network $\mathbf{\epsilon}_{\phi_{\mathrm{est}}}(\mathbf{z}_{t};y,p,c,t)$, which is trained on the rendered images by $\{\theta\}_{i=1}^{k}$. The gradient of VPSD can be formed as,
@@
In [45], only selected particles update the LoRA network in each iteration. However, this approach neglects the learning progression of LoRA networks, which are used to represent variational distributions. These networks typically require numerous iterations to approximate the optimal distribution, resulting in slow convergence. Unfortunately, the randomness introduced by particle initialization can lead to early learning of suboptimal particles, which adversely affects the final convergence result. In VPSD, we introduce a Reward Feedback Learning method, as Fig. 3 illustrates. This method leverages a pre-trained reward model [49] to assign reward scores to samples collected from LoRA model. Then LoRA model subsequently updates from these reweighted samples,
@@
Figure 4: Results of qualitative evaluation. Note: DiffSketcher [48] is originally designed for vector sketch generation. We reimplemented it for generating RGB vector images.
@@ 3.3 Vector Representation Primitives
In addition to text prompts, SVGDreamer provides a variety of vector representations for style control. These vector representations are achieved by limiting primitive types and their parameters. Users can control the art style generated by SVGDreamer by modifying the input text or by constraining the set of primitives and parameters. We explore six settings: 1) Iconography is the most common SVG style, consists of several paths and their fill colors. This style allows for a wide range of compositions while maintaining a minimalistic expression. We utilize closed form Bézier curves with trainable control points and fill colors. 2) Sketch is a way to convey information with minimal expression. We use open form Bézier curves with trainable control points and opacity. 3) Pixel Art is a popular video-game inspired style, frequently used for character and background art. We utilize square SVG polygons with customizable fill colors. The control point of the polygon is a constant value, and the color attribute is trainable. 4) Low-Poly is to consciously cut and pile up a certain number of simple geometric shapes according to the modeling laws of objects. We use square SVG polygons with trainable control points and fill colors. 5) Painting is a means of approximating the painter’s painting style in vector graphics. We use open form Bézier curves with trainable control points, stroke colors and stroke widths. 6) Ink and Wash Painting is a traditional Chinese art form that utilizes varying concentrations of black ink. We use open form Bézier curves with trainable control points, opacity, and stroke widths.
Figure 1 displays a range of SVG styles consisting of six vector primitives. Additional examples can be found in Fig. 8 and Fig. 9.

### 17 Mar 2024 v3

@@ Abstract
Recently, text-guided scalable vector graphics (SVGs) synthesis has shown promise in domains such as iconography and sketch. However, existing text-to-SVG generation methods lack editability and struggle with visual quality and result diversity. To address these limitations, we propose a novel text-guided vector graphics synthesis method called SVGDreamer. SVGDreamer incorporates a semantic-driven image vectorization (SIVE) process that enables the decomposition of synthesis into foreground objects and background, thereby enhancing editability. Specifically, the SIVE process introduce attention-based primitive control and an attention-mask loss function for effective control and manipulation of individual elements. Additionally, we propose a Vectorized Particle-based Score Distillation (VPSD) approach to tackle the challenges of shape over-smoothing, color over-saturation, limited diversity in results, and slow convergence in existing text-to-SVG generation methods. VPSD models SVGs as distributions of control points and colors to counteract over-smoothing and over-saturation. Furthermore, VPSD leverages a reward model to reweight vector particles, which improves aesthetic appeal and accelerates convergence. Extensive experiments have been conducted to validate the effectiveness of SVGDreamer, demonstrating its superiority over baseline methods in terms of editability, visual quality, and diversity. Project page: https://ximinng.github.io/SVGDreamer-project/
@@ 1 Introduction
Scalable Vector Graphics (SVGs) represent visual concepts using geometric primitives such as Bézier curves, polygons, and lines. Due to their inherent nature, SVGs are highly suitable for visual design applications, such as posters and logos. Secondly, compared to raster images, vector images can maintain compact file sizes, making them more efficient for storage and transmission purposes. More importantly, vector images offer greater editability, allowing designers to easily select, modify, and compose elements. This attribute is particularly crucial in the design process, as it allows for seamless adjustments and creative exploration.
@@
In recent years, there has been a growing interest in general vector graphics generation. Various optimization-based methods [4, 29, 20, 40, 41, 34, 12, 48] have been proposed, building upon the differentiable rasterizer DiffVG [14]. These methods, such as CLIPDraw [4] and VectorFusion [12], differ primarily in their approach to supervision. Some works [4, 29, 20, 34, 40, 41] combine the CLIP model [24] with DiffVG [14], using CLIP as a source of supervision. More recently, the significantly progress achieved by Text-to-Image (T2I) diffusion models [21, 27, 25, 28, 37] has inspired the task of text-to-vector-graphics. Both VectorFusion [12] and DiffSketcher [48] attempted to utilize T2I diffusion models for supervision. These models make use of the high-quality raster images generated by T2I models as targets to optimize the parameters of vector images. Additionally, the priors embedded within T2I models can be distilled and applied in this task. Consequently, models that use T2I for supervision generally perform better than those using the CLIP model.
@@
Despite their impressive performance, existing T2I-based methods have certain limitations. Firstly, the vector images generated by these methods lack editability. Unlike the conventional approach of creating vector graphics, where individual elements are added one by one, T2I-based methods do not distinguish between different components during synthesis. As a result, the objects become entangled, making it challenging to edit or modify a single object independently. Secondly, there is still a large room for improvement in visual quality and diversity of the results generated by these methods. Both VectorFusion [12] and DiffSketcher [48] use data augmentation extended Score Distillation Sampling (SDS) [23] to distill priors from the T2I models. However, it has been observed that SDS can lead to issues such as color over-saturation and over-smoothing, resulting in a lack of fine details in the generated vector images. Besides, SDS optimizes a set of control points in the vector graphic space to obtain the average state of the vector graphic corresponding to the text prompt in a mode-seeking manner [23]. This leads to a lack of diversity and detailed construction in the SDS-based approach [12, 48], along with absent text prompt objects.
@@
To address the aforementioned issues, we present a new model called SVGDreamer for text-guided vector graphics generation. Our primary objective is to produce vector graphics of superior quality that offer enhanced editability, visual appeal, and diversity. To ensure editability, we propose a semantic-driven image vectorization (SIVE) process. This approach incorporates an innovative attention-based primitive control strategy, which facilitates the decomposition of the synthesis process into foreground objects and background. To initialize the control points for each foreground object and background, we leverage cross-attention maps queried by text tokens. Furthermore, we introduce an attention-mask loss function, which optimizes the graphic elements hierarchically. The proposed SIVE process ensures the separation and editability of the individual elements, promoting effective control and manipulation of the resulting vector graphics.
@@
To improve the visual quality and diversity of the generated vector graphics, we introduce a new technique: Vectorized Particle-based Score Distillation (VPSD) for vector graphics refinement. Previous works in vector graphics synthesis [12, 48, 11] that utilized score distillation sampling [23] often encountered issues like shape over-smoothing, color over-saturation, limited diversity in results, and slow convergence in the synthesis results [23, 48]. To address these challenges, VPSD models SVGs as distributions of control points and colors, respectively. VPSD adopts a LoRA [10] network to estimate the distributions, aligning vector graphics with the pretrained diffusion model. Furthermore, to enhance the aesthetic appeal of the generated vector graphics, we introduce ReFL [49] to fine-tune the estimation network. Through this refinement process, we achieve the final vector graphics that exhibit high editability, superior visual quality, and increased diversity. To validate the effectiveness of our proposed method, we perform extensive experiments to evaluate the model across multiple aspects. In summary, our contributions can be summarized as follows:
@@
  • We present the semantic-driven image vectorization (SIVE) method, which ensures that the generated vector objects are separate and flexible to edit. Additionally, we propose the vectorized particle-based score distillation (VPSD) loss to guarantee that the generated vector graphics exhibit both exceptional visual quality and a wide range of diversity.
@@ 2.1 Vector Graphics Generation
Scalable Vector Graphics (SVGs) offer a declarative format for visual concepts expressed as primitives. One approach to creating SVG content is to use Sequence-To-Sequence (seq2seq) models to generate SVGs [5, 16, 1, 26, 43, 44, 46]. These methods heavily rely on dataset in vector form, which limits their generalization ability and their capacity to synthesize complex vector graphics. Instead of directly learning an SVG generation network, an alternative method of vector synthesis is to optimize towards a matching image during evaluation time.
Li et al. [14] introduce a differentiable rasterizer that bridges the vector graphics and raster image domains. While image generation methods that traditionally operate over vector graphics require a vector-based dataset, recent work has demonstrated the use of differentiable renderers to overcome this limitation [31, 39, 26, 29, 18, 38, 36, 48]. Furthermore, recent advances in visual text embedding contrastive language-image pre-training model (CLIP) [24] have enabled a number of successful methods for synthesizing sketches, such as CLIPDraw[4], CLIP-CLOP [20], and CLIPasso [40]. A very recent work VectorFusion [12] and DiffSketcher [48] combine differentiable renderer with text-to-image diffusion model for vector graphics generation, resulting in promising results in fields such as iconography, pixel art, and sketch.
@@ 2.2 Text-to-Image Diffusion Model
Denoising diffusion probabilistic models (DDPMs) [32, 33, 8, 35], particularly those conditioned on text, have shown promising results in text-to-image synthesis. For example, Classifier-Free Guidance (CFG) [7] has improved visual quality and is widely used in large-scale text conditional diffusion model frameworks, including GLIDE [21], Stable Diffusion [27], DALL·E 2 [25], Imagen [28] and DeepFloyd IF [37]. The progress achieved by text-to-image diffusion models [21, 27, 25, 28] also promotes the development of a series of text-guided tasks, such as text-to-3D [23]. In this work, we employ Stable Diffusion model to provide supervision for text-to-SVG generation.
@@ 3 Methodology
Figure 2: Overview of SVGDreamer. The method consists of two parts: semantic-driven image vectorization (SIVE) and SVG synthesis through VPSD optimization. The result obtained from SIVE can be used as input for VPSD, which helps in establishing a dependable initialization point.
@@ 3.2 Vectorized Particle-based Score Distillation
While vectorizing a rasterized diffusion sample is lossy, recent techniques [12, 48] have identified the SDS loss [23] as beneficial for our task of generating vector graphics. To synthesize a vector image that matches a given text prompt $y$, they directly optimize the parameters $\theta=\{P_{i},C_{i}\}_{i=1}^{n}$ of a differentiable rasterizer $\mathcal{R}(\theta)$ via SDS loss. At each iteration, the differentiable rasterizer is used to render a raster image $\mathbf{x}=\mathcal{R}(\theta)$, which is augmented to obtain a $\mathbf{x}_{a}$. Then, the pretrained latent diffusion model (LDM) $\epsilon_{\phi}$ uses a VAE encoder [3] to encode $\mathbf{x}_{a}$ into a latent representation $\mathbf{z}=\mathcal{E}(\mathbf{\mathbf{x}_{a}})$, where $\mathbf{z}\in\mathbb{R}^{(H/f)\times(W/f)\times 4}$ and $f$ is the encoder downsample factor. Finally, the gradient of SDS is estimated by,
@@
Unfortunately, SDS-based methods often suffer from issues such as shape over-smoothing, color over-saturation, limited diversity in results, and slow convergence in synthesis results [23, 12, 48, 11].
Figure 3: An illustration of the VPSD. VPSD accepts $k$ sets of SVG parameters as input. The rendered image is sent to the pretrained diffusion and the score of the SVG parameters distribution to compute the gradient of VPSD.
Inspired by the principled variational score distillation framework [45], we propose vectorized particle-based score distillation (VPSD) to address the aforementioned issues. Instead of modeling SVGs as a set of control points and corresponding colors like SDS, we model SVGs as the distributions of control points and colors respectively. In principle, given a text prompt $y$, there exists a probabilistic distribution $\mu$ of all possible vector shapes representations. Under a vector representation parameterized by $\theta$, such a distribution can be modeled as a probabilistic density $\mu(\theta|y)$. Compared with SDS that optimizes for the single $\theta$, VPSD optimizes for the whole distribution $\mu$, from which we can sample $\theta$. Motivated by previous particle-based variational inference methods, we maintain $k$ groups of vector parameters $\{\theta\}_{i=1}^{k}$ as particles to estimate the distribution $\mu$, and $\theta(i)$ will be sampled from the optimal distribution $\mu^{\ast}$ if the optimization converges. This optimization can be realized through two score functions: one that approximates the optimal distribution with a noisy real image, and one that represents the current distribution with a noisy rendered image. The score function of noisy real images can be approximated by the pretrained diffusion model [27] $\mathbf{\epsilon}_{\phi}(\mathbf{z}_{t};y,t)$. The score function of noisy rendered images is estimated by another noise prediction network $\mathbf{\epsilon}_{\phi_{\mathrm{est}}}(\mathbf{z}_{t};y,p,c,t)$, which is trained on the rendered images by $\{\theta\}_{i=1}^{k}$. The gradient of VPSD can be formed as,
@@
In [45], only selected particles update the LoRA network in each iteration. However, this approach neglects the learning progression of LoRA networks, which are used to represent variational distributions. These networks typically require numerous iterations to approximate the optimal distribution, resulting in slow convergence. Unfortunately, the randomness introduced by particle initialization can lead to early learning of suboptimal particles, which adversely affects the final convergence result. In VPSD, we introduce a Reward Feedback Learning method, as Fig. 3 illustrates. This method the significantrages a pre-trained reward model [49] to assign reward scores to samples collected from LoRA model. Then LoRA model subsequently updates from these reweighted samples,
@@
Figure 4: Results of qualitative evaluation. Note: DiffSketcher is originally designed for vector sketch generation. We reimplemented it for generating RGB vector images.
@@ 3.3 Vector Representation Primitives
In addition to text prompts, SVGDreamer provides a variety of vector representations for style control. These vector representations are achieved by limiting primitive types and their parameters. Users can control the art style generated by SVGDreamer by modifying the input text or by constraining the set of primitives and parameters. We explore six settings: 1) Iconography is the most common SVG style, consists of several paths and their fill colors. This style allows for a wide range of compositions while maintaining a minimalistic expression. We utilize closed form Bézier curves with trainable control points and fill colors. 2) Sketch is a way to convey information with minimal expression. We use open form Bézier curves with trainable control points and opacity. 3) Pixel Art is a popular video-game inspired style, frequently used for character and background art. We use square SVG polygons with fill colors. 4) Low-Poly is to consciously cut and pile up a certain number of simple geometric shapes according to the modeling laws of objects. We use square SVG polygons with trainable control points and fill colors. 5) Painting is a means of approximating the painter’s painting style in vector graphics. We use open form Bézier curves with trainable control points, stroke colors and stroke widths. 6) Ink and Wash Painting is a traditional Chinese art form that utilizes varying concentrations of black ink. We use open form Bézier curves with trainable control points, opacity, and stroke widths.